OpenStudy (anonymous):

f(x) = x(lnx)^2 find the local max and mins

OpenStudy (anonymous):

I think you find the first derivative and then find the zeros.

OpenStudy (anonymous):

First derivative set equal to zero fr the crit. points. Plug them into the second derivative to determine if they are a max or min.

OpenStudy (anonymous):

Yess i got up to there. But i cannot solve 0=(lnx)(2x+lnx)

OpenStudy (anonymous):

how do you know that?

OpenStudy (anonymous):

You took the derivative incorrectly. You need to use the product and chain rule here.

OpenStudy (anonymous):

suppoed to be lnx(2+lnx)

OpenStudy (anonymous):

but how can you solve lnx=0?

OpenStudy (anonymous):

ln(1)=0

OpenStudy (anonymous):

in order to cancel the ln you have to e^

OpenStudy (anonymous):

because of that you also have to do it to the right side. e^(lnx) = e^(0) x = 1 anything to the zero is 1.